OpenStudy (anonymous):

Solve: 3(x + 1) + 6 = 33

OpenStudy (anonymous):

3x + 3 = 33-6 3x = 33-6-3 3x = 24 x = 24/3 x = 8

OpenStudy (anonymous):

thanx

OpenStudy (anonymous):

3x+3 + 6 = 33 3x = 33-6-3 solve it

OpenStudy (anonymous):

you're welcome :)